A leading magazine (like Barrons) reported at one time that the average number of weeks an individual is unemployed is 28 weeks. Assume that for the population of all unemployed individuals the population mean length of unemployment is 28 weeks and that the population standard deviation is 7 weeks Suppose you would like to select a random sample of 40 unemployed individuals for a follow-up study Find the probability that a single randomly selected value is less than 29 PX29) Find the probability that a sample of size n = 40 is randomly selected with a mean less than 29. P(M<29)- Enter your answers as numbers accurate to 4 decimal places. License Points possible: 1 Unlimited attempts Submit
Answer #1

Solution:

Given that,

mean = \mu = 28

standard deviation = \sigma = 7

a ) P ( x  <  29 )

=p ( x - \mu/\sigma)  \leq (29 - 28 / 7 )

= p( z < 1 / 7 )

= p ( z < 0.14 )   

Using z table

= 0.5557

Probability = 0.5557

b ) n = 40

So,

\bar x = M

  \mu\bar x = 28

\sigma\bar x =  (\sigma /\sqrtn) = (7 / \sqrt40 ) = 1.1068

P ( M <  29 )

=p ( M - \mu/\sigma)  \leq (29 - 28 / 1.1068 )

= p( z < 1 / 1.1068 )

= p ( z < 0.90 )   

Using z table

= 0.8159

Probability = 0.8159